OpenStudy (jennychan12):

The derivative of f is given f'(x) = e^x(-x^3+3x)-3 for [0,5]. What value of f(x) is an absolute minimum?

OpenStudy (jennychan12):

\[f'(x) = e^x(-x^2+3x)-3 \] for \[0 \le x \le 5\]

OpenStudy (jennychan12):

i used a graphing calculator to find the x-coordinate at the minimum value. is there a way to do this without a graphing calculator?

OpenStudy (anonymous):

The most you could do without a graphing calculator is set \(f'(x)=0\) and try solving for \(x\), as you normally would with the first derivative test. But solving for \(x\) by hand would be very, very difficult.

OpenStudy (jennychan12):

yeah, i tried solving for x. it's kinda frustrating. and annoying. and this question is part of the no calculator, so i dunno how i would do it. :(

OpenStudy (jennychan12):

thanks anyways :)